OpenStudy (acxbox22):

List the transformations from the graph of y=1/x to y=1/6x

OpenStudy (acxbox22):

@mathmale

OpenStudy (acxbox22):

@hero @jim_thompson5910

OpenStudy (acxbox22):

@iambatman

jimthompson5910 (jim_thompson5910):

the equation is \[\Large y=\frac{1}{6x}\] right?

OpenStudy (acxbox22):

yes

OpenStudy (acxbox22):

is it a vertical stretch by a factor of 1/6?

jimthompson5910 (jim_thompson5910):

good, since \[\Large y=\frac{1}{6x} = \frac{1}{6}*\frac{1}{x}\]

jimthompson5910 (jim_thompson5910):

so it's 1/6 as high as it used to be

OpenStudy (acxbox22):

vertical shrink right?